# Steering Large Language Models using Conceptors: Improving Addition-Based Activation Engineering

### v1

@@ Abstract
Large language models have transformed AI, yet reliably controlling their outputs remains a challenge. This paper explores activation engineering, where outputs of pre-trained LLMs are controlled by manipulating their activations at inference time. Unlike traditional methods using a single steering vector, we introduce conceptors–mathematical constructs that represent sets of activation vectors as ellipsoidal regions. Conceptors act as soft projection matrices and offer more precise control over complex activation patterns. Our experiments demonstrate that conceptors outperform traditional methods across multiple in-context learning steering tasks. We further use Boolean operations on conceptors that allows for combined steering goals that empirically outperforms combining steering vectors on a set of tasks. These results highlight conceptors as a promising tool for more effective steering of LLMs.
@@ 1 Introduction
Large language models (LLMs) have rapidly advanced AI capabilities [1], but their potential to spread misinformation [2], reinforce biases [3], and develop harmful behaviors [4] highlights the urgent need for methods to understand and control their outputs. Various methods, including reinforcement learning from human feedback (RLHF) [5], supervised fine-tuning [6], and prompt engineering [7], have been proposed to steer LLM outputs toward desired patterns. However, RLHF and fine-tuning are computationally expensive and struggle with generalization [8, 9], while prompt engineering often produces inconsistent results [10]. Activation engineering [11, 12] has recently been proposed as a new steering method which works by directly modifying the model’s activations at inference time without changing the model’s parameters and without expensive optimization. A steering vector that represents desired behavior can be computed directly or (more commonly) contrastively from positive and negative example [13]. However, finding contrastive prompts to demonstrate complex patterns is not always possible, and most importantly, the performance of activation addition is not always reliable [12]. This paper introduces an alternative to the current predominant approach for steering LLMs using activation engineering. Instead of averaging or subtracting a set of activation vectors to form a steering vector, the cached activations are used to compute a conceptor, which we refer to as a “steering matrix”. Instead of manipulating the LLM’s activations using vector addition, the activations are (softly) projected using a matrix-vector multiplication with the steering matrix. We contribute the following: (1) we introduce a novel conceptor steering mechanism for LLMs, (2) we apply this mechanism to function vectors on GPT-NeoX and GPT-J, and (3) we show how a Boolean algebra on conceptors can be used for combining functions on GPT-J.
@@ 2 Background
Adding steering vectors to the residual stream has successfully been used to control the output of LLMs across various domains [12, 13, 14]. The use case that will mainly be focused on here are the findings from the paper by Todd et al. [15]. Their work showed that a steering vector can be extracted from the residual stream that captures the activation space of an input-output function (e.g. a function that takes a word and returns its antonym). This steering vector can then be added to the residual stream at inference time to steer the model toward performing the captured function.
Their baseline method works as follows. First, a set of in-context learning (ICL) prompts $P_{f}$ that demonstrate a particular task $f$ (the execution of an input-output function) are compiled. Then for each prompt $p_{i}^{f}\in P_{f}$ (e.g., $p_{1}^{antonym}=$ hot:cold,old:) the final token’s activations $h_{\ell}(p_{i}^{f})$ are cached at a specific layer $\ell$ from the residual stream $h$. The cached activation vectors are then averaged into the steering vector $\bar{h}_{\ell}^{f}$ for task $f$ at layer $\ell$:
@@
Figure 1: Illustration showing the basic geometric difference between additive and conceptor steering using a set of activations for the antonym task.
@@ 3 Conceptors as Steering Matrices
Conceptors can broadly be defined as a neuro-computational mechanism designed to encapsulate and manipulate the state space of neural activations [17]. A conceptor matrix $C$ is a positive semi-definite matrix that captures the principal directions and variances of a set of neural activation vectors. This structure can be visualized as a high-dimensional ellipsoid that describes the overall shape and spread of the activations’ “underlying pattern”, or state space region. Because conceptors are computed from the cloud of activation vectors and encode the correlations between activations, conceptors can better capture the activation space of complex patterns compared to simple point representations, which discard information about correlations.
@@
where $X$ is a matrix of neural activation vectors (stacked as rows), $\|\cdot\|_{F}$ is the Frobenius norm, and $\alpha$ is the regularization parameter also referred to as the conceptor’s aperture. This aperture parameter $\alpha$ balances the trade-off between accurately representing the activation pattern and maintaining a generalized representation. When $\alpha$ is large, the eigenvalues $\mu_{i}$ approach 1 and $C$ approaches the identity matrix, causing the conceptor to allow for more signal components to pass through the projection of the states with the conceptor matrix $Cx$. Conversely, when $\alpha$ is small, the eigenvalues $\mu_{i}$ approach 0, causing the conceptor to allow for less variability. In the extreme case of $\alpha=0$, the conceptor collapses to the zero mapping.
The closed-form solution to this problem is given by:
$C(R,\alpha)=R\left(R+\alpha^{-2}I\right)^{-1}\quad\text{where}\quad R=\frac{X^{T}X}{n}$ (3)
@@
We can use the conceptor for steering by collecting activations $h_{\ell}(p_{i}^{f})$ into $X$ and then compute the associated conceptor $C_{\ell}^{f}$ using Equation 3, and then steer new hidden activations with:
@@
We can combine multiple steering matrices using the conceptor Boolean operations as defined by Jaeger [18]. For our experiments, we use the AND operation which combines two conceptors $C_{1}$ and $C_{2}$ into a joint conceptor $C_{1}\wedge C_{2}=\left(C_{1}^{-1}+C_{2}^{-1}-I\right)^{-1}$. See Appendix for A.1.1 for more details.
@@ 4 Experiments
For our experiments, we will use EleutherAI’s GPT-J 6B and GPT-NeoX 20B models, as done in previous works [15, 16]. For all experiments, we find optimal hyperparameters for each steering method at every layer. The details of our grid search for $\alpha$ and $\beta_{c}$ for conceptor-based steering and $\beta_{\text{add}}$ for additive steering can be found in Appendix A.2.2.
@@ 4.1 Function Steering
The results in Figure 2 show that conceptor-based steering outperforms additive steering (the baseline method reported in Ref. [15]) for every task on both tested models. The performance improvement of the proposed conceptor mechanism is as large as In line with previous findings [15, 16], steering is most effective across layers 9-16 for GPT-J and layers 10-30 for GPT-NeoX.
@@ 4.2 Steering Composite Functions
We further conducted experiments where two conceptors representing three different compound functions were combined using the AND operator. The input-output example dataset for this function was generated using GPT-4o.
Figure 3: Performance of additive steering and conceptor steering on composite functions. For explanation of the figure caption, see text.
To present the baseline for how well non-combined steering mechanisms perform, we show results for the conceptor $C^{1,2}$ and the steering vector $\bar{h_{\ell}^{1,2}}$ that were each computed on the compound function directly. We then combine the conceptors computed on the individual functions $C^{1}$ and $C^{2}$ using the AND operation as $C^{1}\wedge C^{2}$, and we combine the steering vectors $\bar{h_{\ell}^{1}}$ and $\bar{h_{\ell}^{2}}$ using their arithmetic mean $\frac{1}{2}(\bar{h_{\ell}^{1}}+\bar{h_{\ell}^{2}})$. Figure 3 shows the performance of all compared methods across all layers of the GPT-J model. In line with the results from Section 4.1, the conceptor baseline outperformed the additive baseline on all three tasks. The AND-combined conceptor outperformed the mean-combined steering vectors. On one of the three tasks, english-french & antonyms, the AND-combined conceptor even outperforms the additive baseline.
@@ 5 Conclusion
In our experiments, conceptor-based steering generally outperformed addition-based methods. That said, further research should be conducted to assess the mechanisms’ impact on the model’s overall capabilities, as well as the performance on more complex behaviors/tasks. Although we tested the steering performance on two models, examining scalability on even larger models, such as LLaMA-70B, could provide valuable insights.
A limitation of conceptors is their reliance on more data points to build accurate representations. In addition to this, the inherent mathematical structure and additional required computations makes it relatively more computationally expensive compared to simple addition-based methods. Conceptors also introduce a new hyperparameter, aperture, that requires tuning for optimal performance.
Despite these challenges, conceptor-based steering methods could offer a more precise and effective way to steer LLMs compared to traditional addition-based methods, proposing a fundamental shift in what is possible with activation engineering. While acknowledging their potential, further research is needed to fully assess their broader applicability and limitations.
@@ Appendix A Appendix
A.1 Conceptors
Conceptors are mathematical constructs that can be used for the management of neural activations [17]. A conceptor can be visualized as a structure that describes the activational pattern, or state cloud, of a set of high-dimensional activation points using an ellipsoid (see Figure 4). This conceptor is mathematically represented by a positive semi-definite matrix with eigenvalues between zero and unity that can be used to (softly) project a new set of activations toward the described ellipsoid.
Figure 4: 2D visualization of 3 Conceptors that describe the "underlying pattern" or state space region of 3 different sets of neural activations.
The eigenvalues $\mu_{i}$ of the conceptor matrix $C$ are defined as:
where $\lambda_{i}$ represents the eigenvalues of the correlation matrix $R$. These eigenvalues $\mu_{i}$ fall within the interval $[0,1]$ and are influenced by the aperture parameter $\alpha$. When $\alpha$ is large, the eigenvalues $\mu_{i}$ approach 1 and $C$ approaches the identity matrix, causing the conceptor to allow for more signal components to pass through the projection of the states with the conceptor matrix $Cx$. Conversely, when $\alpha$ is small, the eigenvalues $\mu_{i}$ approach 0, causing the conceptor to allow for less variability. In the extreme case of $\alpha=0$, the conceptor collapses to the zero mapping.
A.1.1 Boolean Operations on Conceptors
In this section, we introduce the Boolean operations on conceptors that were originally defined by Jaeger [18]. We begin with the OR operation on conceptors, which can be interpreted as merging the data from which each conceptor is computed by adding the covariance matrices on which $C_{1}$ and $C_{2}$ were computed. Given that $C_{1}$ was computed with the covariance matrix $R_{1}$ and $C_{2}$ was computed with the covariance matrix $R_{2}$, the conceptor that is computed on the sum of the two covariance matrices $R_{1}+R_{2}$ is defined as $C_{1}\vee C_{2}$:
$C_{1}\vee C_{2}$ $=(R_{1}+R_{2})(R_{1}+R_{2}+\alpha^{-2}I)^{-1}$ (5)
$C_{1}\vee C_{2}$ $=\left(I+\left(C(I-C)^{-1})+B(I-B)^{-1}\right)^{-1}\right)^{-1}$ (6)
The NOT operation on a conceptor $C$ is defined as the conceptor $\neg C$ that is computed on a covariance matrix $R^{-1}$ that is inverse to the original covariance matrix $R$ for conceptor $C$. Intuitively, $\neg C$ can be interpreted as the conceptor that arises from data that which co-vary inversely to the data giving rise to $C$:
$\neg C$ $=R^{-1}(R^{-1}+\alpha^{-2}I)^{-1}$ (7)
$\neg C$ $=I-C$ (8)
The AND operation can now be obtained using de Morgan’s law $a\wedge b=\not(a\vee b)$ such that the conceptor $C_{1}\wedge C_{2}$ is computed using the correlation matrix $(R_{1}^{-1}+R_{2}^{-1})^{-1}$. This leads to:
$C_{1}\wedge C_{2}$ $=(C_{1}^{-1}+C_{2}^{-1}+I)^{-1}$ (10)
A.2 Experimental Details
@@ A.2.2 Hyperparameter optimization
A.3 Additional Experimental Results
A.3.1 Mean centering
@@
antonyms capitalize country-capital english-french present-past
Addition 20.54% 93.16% 32.04% 18.88% 69.66%
Addition (MC) 31.20% 95.00% 63.90% 34.32% 83.32%
Conceptor 52.14% 96.68% 81.62% 59.02% 91.56%
Conceptor (MC) 52.82% 96.26% 85.32% 61.32% 91.88%
Table 1 and Figure 6 show that mean-centering provides a small improvement for both addition-based and conceptor-based steering. The experimental setup is as described in Appendix A.2.1. Indeed, mean-centering improves the performance of additive steering by as much as 99% (on the country-capital task). For conceptor-based steering the improvements of mean-centering are relatively smaller – at most 5% on the country-capital task. Conceptor-based steering outperforms additive steering on all tasks, even comparing additive steering with mean-centering against conceptor-based steering without mean-centering.
Figure 6: The effect of mean centering on conceptor-based and addition-based steering on the GPT-J (6B) model across all layers, computed on five different function vector tasks. The line shows the best average performance across five runs for the best hyperparameters for the given layer.
A.4 Hyperparameter Sweep Results

### v4

@@ Abstract
Large language models have transformed AI, yet reliably controlling their outputs remains a challenge. This paper explores activation engineering, where outputs of pre-trained LLMs are controlled by manipulating their activations at inference time. Unlike traditional methods using a single steering vector, we introduce conceptors–mathematical constructs that represent sets of activation vectors as ellipsoidal regions. Conceptors act as soft projection matrices and offer more precise control over complex activation patterns. Our experiments demonstrate that conceptors outperform traditional methods across multiple steering tasks. We further use Boolean operations on conceptors for combined steering goals that empirically outperform additively combining steering vectors on a set of tasks. These results highlight conceptors as a promising tool for more effective steering of LLMs. Our code is available on github.com/jorispos/conceptorsteering.
@@ 1 Introduction
Large language models (LLMs) have rapidly advanced AI capabilities [1], but their potential to spread misinformation [2], reinforce biases [3], and develop harmful behaviors [4] highlights the urgent need for methods to understand and control their outputs. Various methods, including reinforcement learning from human feedback (RLHF) [5], supervised fine-tuning [6], and prompt engineering [7], have been proposed to steer LLM outputs toward desired patterns. However, RLHF and fine-tuning are computationally expensive and struggle with generalization [8, 9], while prompt engineering often produces inconsistent results [10].
Activation engineering [11, 12] has recently been proposed as a new steering method which works by directly modifying the model’s activations at inference time without changing the model’s parameters and without expensive optimization. A steering vector that represents desired behavior can be computed directly or (more commonly) contrastively from positive and negative examples [13]. However, finding contrastive prompts to identify complex patterns is not always possible and, more importantly, the performance of activation addition for steering is not reliable [12].
This paper introduces an alternative to the predominant approach for steering LLMs using activation engineering. Instead of averaging or subtracting a set of activation vectors to form a steering vector, we use the cached activations to compute a conceptor [14], which we refer to as a “steering matrix”. Instead of manipulating the LLM’s activations using vector addition, the activations are (softly) projected using a matrix-vector multiplication with the steering matrix. We contribute the following: (1) we introduce a novel application of conceptors [14] as steering mechanisms for LLMs, (2) we apply this mechanism to function vectors [15] on GPT-NeoX and GPT-J, and (3) we show how a Boolean algebra on conceptors [16] can be used for combining steering targets on GPT-J.
@@ 2 Background
Adding steering vectors to the residual stream has been used to control the output of LLMs across various domains [12, 13, 17]. The use case that will mainly be focused on here are the findings from the paper by Todd et al. [15]. Their work showed that a steering vector can be extracted from the residual stream that captures the activation space of an input-output function (e.g. a function that takes a word and returns its antonym). This steering vector can then be added to the residual stream at inference time to steer the model toward performing the captured function. See Figure 6(b) in Appendix A.1.1 for an illustration of function vector tasks.
Figure 1: Illustration showing the basic geometric difference between additive and conceptor steering using a set of activations for the antonym task. Additive steering acts as a translation of the activation vectors by a fixed steering vector. Conceptor steering acts as a (soft) projection onto a target ellipsoid.
Their baseline method works as follows. First, a set of in-context learning (ICL) prompts $P_{f}$ that demonstrate a particular task $f$ (the execution of an input-output function) are compiled. Then, for each prompt $p_{i}^{f}\in P_{f}$ (e.g., $p_{1}^{antonym}=$ hot:cold,old:), the final token’s activations $h_{\ell}(p_{i}^{f})$ are cached at a specific layer $\ell$ from the residual stream $h$. The cached activation vectors are then averaged into the steering vector $\bar{h}_{\ell}^{f}$ for task $f$ at layer $\ell$:
@@ 3 Conceptors as Steering Matrices
Conceptors can broadly be defined as a neuro-computational mechanism designed to encapsulate and manipulate the state space of neural activations [14]. A conceptor matrix $C$ is a positive semi-definite matrix that captures the principal directions and variances of a set of neural activation vectors. This structure can be visualized as a high-dimensional ellipsoid that describes the overall shape and spread of the activations’ “underlying pattern”, or state space region. See Figure 2 for a visual illustration. Because conceptors are computed from the cloud of activation vectors and encode the correlations between activations, conceptors can better capture the activation space of complex patterns compared to simple point representations, which discard information about correlations. This difference is illustrated geometrically in Figure 1.
Figure 2: Illustration of three conceptors as ellipsoids that capture the state space region of different sets of neural activations in 3D space (black points). Reproduced from Jaeger [14].
@@
where $X$ is a matrix of neural activation vectors (stacked as rows), $\|\cdot\|_{F}$ is the Frobenius norm, and $\alpha$ is the regularization parameter also referred to as the conceptor’s aperture. This aperture parameter $\alpha$ balances the trade-off between accurately representing the activation pattern and maintaining a generalized representation. The closed-form solution to this problem is given by:
$C(R,\alpha)=R\left(R+\alpha^{-2}I\right)^{-1}\quad\text{with}\quad R=\frac{X^{T}X}{n}$ (3)
@@
The eigenvalues $\mu_{i}$ of the conceptor matrix $C$ are defined as:
where $\lambda_{i}$ represents the eigenvalues of the correlation matrix $R$. These eigenvalues $\mu_{i}$ fall within the interval $[0,1]$ and are influenced by the aperture parameter $\alpha$. When $\alpha$ is large, the eigenvalues $\mu_{i}$ approach 1 and $C$ approaches the identity matrix, causing the conceptor to allow for more signal components to pass through the projection of the states with the conceptor matrix $Cx$. Conversely, when $\alpha$ is small, the eigenvalues $\mu_{i}$ approach 0, causing the conceptor to allow for less variability. In the extreme case of $\alpha\rightarrow 0$, the conceptor tends to the zero mapping.
We can use the conceptor for steering by collecting activations $h_{\ell}(p_{i}^{f})$ into $X$ and then compute the associated conceptor $C_{\ell}^{f}$ using Equation 3, and finally steer new hidden activations $h_{\ell}$ with:
@@
3.1 Boolean Operations on Conceptors
We can combine multiple steering matrices using the conceptor Boolean operations as defined by Jaeger [16]. We begin with the OR operation on conceptors, which can be interpreted as merging the data from which each conceptor is computed by adding the covariance matrices on which $C_{1}$ and $C_{2}$ were computed. Given that $C_{1}$ was computed with the covariance matrix $R_{1}$ and $C_{2}$ was computed with the covariance matrix $R_{2}$, the conceptor that is computed on the sum of the two covariance matrices $R_{1}+R_{2}$ is defined as $C_{1}\vee C_{2}$:
$C_{1}\vee C_{2}$ $=(R_{1}+R_{2})(R_{1}+R_{2}+\alpha^{-2}I)^{-1}$ (5)
$C_{1}\vee C_{2}$ $=\left(I+\left(C_{1}(I-C_{1})^{-1}+C_{2}(I-C_{2})^{-1}\right)^{-1}\right)^{-1}$ (6)
The NOT operation on a conceptor $C$ is defined as the conceptor $\neg C$ that is computed on a covariance matrix $R^{-1}$ that is the inverse of the original covariance matrix $R$ for conceptor $C$. Intuitively, $\neg C$ can be interpreted as the conceptor that arises from data that which co-vary inversely to the data giving rise to $C$:
$\neg C$ $=R^{-1}(R^{-1}+\alpha^{-2}I)^{-1}$ (7)
$\neg C$ $=I-C$ (8)
For our experiments, we use the AND operation which can now be obtained from the NOT and OR operations using de Morgan’s law $a\wedge b=\neg(a\vee b)$ such that the conceptor $C_{1}\wedge C_{2}$ is computed using the correlation matrix $(R_{1}^{-1}+R_{2}^{-1})^{-1}$. This leads to:
$C_{1}\wedge C_{2}$ $=(C_{1}^{-1}+C_{2}^{-1}-I)^{-1}$ (10)
3.2 Computational Complexity of Conceptor Steering
The cost of computing a conceptor steering matrix is dominated by the matrix inversion and matrix-matrix multiplication of the activation correlation matrix $R=XX^{T}/n$ (see Equation 3). This correlation matrix is a $n\times n$-dimensional matrix where $n$ is the dimension of the activation vectors (typically <4096 for the model sizes we presented, or up to 8192 for larger models such as Llama-2-70B), so the complexity of the conceptor computation is $\mathcal{O}(n^{3})$. This computation is done entirely offline and the cost is amortized over all future applications of the steering method. The final conceptor $C\in\mathbb{R}^{n\times n}$ takes $O(n^{2})$ memory – the same amount as a weight matrix acting on the activation vectors. For 32-bit floating point numbers, this amounts to 17MB for $n=2048$, 67MB for $n=4096$, or 268MB for $n=8192$.
During inference, conceptor steering adds an extra matrix-vector multiplication $Cx$ with the activation vector $x$. However, the additional memory and inference cost for applying the conceptor can be eliminated by fusing the conceptor with the succeeding weight matrices for the query, key and value weight matrices. This is equivalent to replacing the existing weight matrix $W_{x}$ with the conceptor-fused weight matrix $W_{x}^{C}=W_{x}C$. This fusing of operations is standard practice when optimizing networks for inference. We note that there may be an overhead cost for switching the conceptor steering on and off which amounts to the cost of changing the network’s computational graph during inference. We believe this overhead to be negligible during auto-regressive generation on a single data sample, but it must be considered when using batch sizes larger than one.
@@ 4 Experiments
For our experiments, we will use EleutherAI’s GPT-J 6B and GPT-NeoX 20B models, as done in previous works on activation steering [15, 18]. For all experiments, we find optimal hyperparameters for each steering method at every layer. The details of our grid search for $\alpha$ and $\beta_{c}$ for conceptor-based steering and $\beta_{\text{add}}$ for additive steering can be found in Appendix A.1.2.
@@ 4.1 Function Steering
The results in Figure 3 show that conceptor-based steering outperforms additive steering (the baseline method reported in Ref. [15]) for every task on both tested models. In line with previous findings [15, 18], steering is most effective across layers 9-16 for GPT-J and layers 10-30 for GPT-NeoX.
Table 1 and Figure 4 show that mean-centering (as outlined in Appendix A.2.1) provides a small improvement for both addition-based and conceptor-based steering. Mean-centering improves the performance of additive steering by as much as 2x (on the country-capital task). For conceptor-based steering the improvements of mean-centering are relatively smaller – at most 5% on the country-capital task. Conceptor-based steering outperforms additive steering on all tasks, even comparing additive steering with mean-centering against conceptor-based steering without mean-centering.
antonyms capitalize country-capital english-french present-past
Addition 20.54% 93.16% 32.04% 18.88% 69.66%
Addition (MC) 31.20% 95.00% 63.90% 34.32% 83.32%
Conceptor 52.14% 96.68% 81.62% 59.02% 91.56%
Conceptor (MC) 52.82% 96.26% 85.32% 61.32% 91.88%
Figure 4: The effect of mean centering on conceptor-based and addition-based steering on the GPT-J (6B) model across all layers, computed on five different function vector tasks (accuracy). The line shows the best average performance across five runs for the best hyperparameters for the given layer.
@@ 4.2 Steering Composite Functions
We further conducted experiments where two conceptors, each representing one of three different functions, were combined using the AND operator. The input-output example dataset for this function was generated using GPT-4o. To present the baseline for how well non-combined steering mechanisms perform, we show results for the conceptor $C^{1,2}$ and the steering vector $\bar{h_{\ell}^{1,2}}$ that were each computed on the compound function directly. We then combine the conceptors computed on the individual functions $C^{1}$ and $C^{2}$ using the AND operation as $C^{1}\wedge C^{2}$, and we combine the steering vectors $\bar{h_{\ell}^{1}}$ and $\bar{h_{\ell}^{2}}$ using their arithmetic mean $\frac{1}{2}(\bar{h_{\ell}^{1}}+\bar{h_{\ell}^{2}})$.
Figure 5: Performance of additive steering and conceptor steering on composite functions. For explanation of the figure caption, see text. Dashed lines represent the “baseline” where the steering mechanism is computed on the composite task. Solid lines show task arithmetic.
Figure 5 shows the performance of all compared methods across all layers of the GPT-J model. In line with the results from Section 4.1, the conceptor baseline outperformed the additive baseline on all three tasks. The AND-combined conceptor outperformed the mean-combined steering vectors. On one of the three tasks, english-french & antonyms, the AND-combined conceptor even outperforms the additive baseline.
@@ 5 Conclusion
In our experiments, conceptor-based steering generally outperformed addition-based methods. Further research should be conducted to assess the mechanisms’ impact on the model’s overall capabilities, the performance on more complex behaviors/tasks, and the scalability to larger models.
A limitation of conceptors is their reliance on more data points to build accurate representations. Additionally, the inherent mathematical structure and additional required computations makes it more computationally expensive compared to simple addition-based methods. However, while more expensive than addition-based approaches, they are still much cheaper than alternatives like RLHF and fine-tuning. Conceptors also introduce a new hyperparameter, the aperture $\alpha$, that may require tuning for optimal performance. In our experiments, we found a single aperture value, $\alpha=0.1$, yields the best performance across all experiments [^1] More precisely, the aperture value $\alpha=0.1$ is within $10\%$ of the best-performing aperture value across all experiments and models. In most experiments, it is the single best value. See Appendix A.3 for more details., but this finding must be verified for new models and steering tasks.
Despite these challenges, conceptor-based steering methods could offer a more precise and effective way to steer LLMs compared to traditional addition-based methods, proposing a fundamental shift in what is possible with activation engineering. Our experiments on conceptor-based steering further suggest that region-based representations may allow for more flexible and nuanced steering compared to point-based representations. The proposed method could have significant positive implications for debiasing models, aligning models with human values, and overall AI safety.
@@ Appendix A Appendix
A.1 Experimental Details
@@ A.1.2 Hyperparameter optimization
A.2 Additional Experimental Results
A.2.1 Mean centering
@@
A.3 Hyperparameter Sweep Results